# EnterpriseGym Corecraft: Training Generalizable Agents on High-Fidelity RL Environments

### v1

@@ Abstract
We show that training AI agents on high-fidelity reinforcement learning environments produces capabilities that generalize beyond the training distribution. We introduce Corecraft, the first environment in EnterpriseGym, Surge AI’s suite of agentic RL environments. Corecraft is a fully operational enterprise simulation of a customer support organization, comprising over 2,500 entities across 14 entity types with 23 unique tools, designed to measure whether AI agents can perform the multi-step, domain-specific work that real jobs demand. Frontier models such as GPT-5.2 and Claude Opus 4.6 solve fewer than 30% of tasks when all expert-authored rubric criteria must be satisfied. Using this environment, we train GLM 4.6 with Group Relative Policy Optimization (GRPO) and adaptive clipping. After a single epoch of training, the model improves from 25.37% to 36.76% task pass rate on held-out evaluation tasks. More importantly, these gains transfer to out-of-distribution benchmarks: +4.5% on BFCL Parallel, +7.4% on $\tau^{2}$-Bench Retail, and +6.8% on Toolathlon (Pass@1). We believe three environment properties are consistent with the observed transfer: task-centric world building that optimizes for diverse, challenging tasks; expert-authored rubrics enabling reliable reward computation; and enterprise workflows that reflect realistic professional patterns. Our results suggest that environment quality, diversity, and realism are key factors enabling generalizable agent capabilities.
@@ 1 Introduction
  • We show transfer to out-of-distribution benchmarks: +4.5% on BFCL Parallel function calling, +7.4% on $\tau^{2}$-Bench Retail customer service, and +6.2% on Toolathlon long-horizon tool use.
@@ 2.1 Agent Benchmarks and Evaluation
Customer service evaluation has been addressed by $\tau$-bench (Yao et al., 2024) and its successor $\tau^{2}$-bench (Barres et al., 2025), which introduced the passk metric measuring reliability across multiple attempts. Toolathlon (Li et al., 2025) benchmarks agents on 108 diverse, long-horizon tasks spanning 32 software applications and 604 tools, with execution-based evaluation. AgentBoard (Ma et al., 2024) provides multi-turn evaluation with subgoal progress tracking across diverse agentic tasks.
@@ 4 Frontier Model Performance on Corecraft
To characterize the difficulty of Corecraft and establish the headroom available for training, we evaluated five frontier models on the current expanded task set, which we refer to as Corecraft-Expanded. This set includes additional challenging tasks added after the training experiments were conducted. Table 1 reports task pass rates, where a task is counted as passed only when all rubric criteria are satisfied across trajectories.
@@
Table 1: Frontier model performance on Corecraft-Expanded, the current expanded task set. Pass rate requires all rubric criteria to be satisfied. Even the best model (GPT-5.2 High) solves fewer than 30% of tasks.
These results reveal substantial headroom. Even GPT-5.2, the strongest model on this set, achieves only 18.18% when all rubric criteria must be satisfied. Most other frontier models fall below 6%, and Claude Sonnet 4.5 does not fully solve any task in Corecraft-Expanded. This confirms that Corecraft presents challenging tasks that expose capability limitations even in the most advanced models, providing a rich training signal for RL.
@@ 5.1 Base Model and Data
We use GLM 4.6 as our base model, a 357B parameter mixture-of-experts architecture with 32B active parameters. The training set comprises tasks from the Corecraft environment, with a held-out evaluation set to measure improvements.
@@ 5.4 Reward Computation
where $C$ is the set of rubric criteria. This provides a dense reward signal while maintaining interpretable task-level metrics. A task is considered passed only when all criteria are satisfied ($r=1$).
@@ 6.1 In-Distribution Performance
Table 3 presents task pass rates on the held-out Corecraft evaluation set used during training. [^1] Since training was conducted, additional tasks have been added to Corecraft, resulting in the Corecraft-Expanded set reported in Section 4. The results in Table 3 reflect the task split available at training time. A task is counted as passed only when all rubric criteria are satisfied, making this a strict measure of end-to-end task completion. After one epoch of training, GLM 4.6 improves from 25.37% to 36.76%, a gain of 11.39 percentage points.
@@ 6.3 Out-of-Distribution Generalization
Table 4 presents results on benchmarks outside the Corecraft training distribution. The trained model shows consistent improvements across both function calling and customer service domains.
@@
Toolathlon (Li et al., 2025) evaluates language agents on 108 diverse, long-horizon tasks spanning 32 software applications and 604 tools, requiring an average of approximately 20 interaction turns per task. The trained model improves from 18.8% to 25.0% (+6.2%), placing it between Claude Haiku 4.5 and Grok-4 in our evaluations. This is a particularly notable transfer result because Toolathlon tasks (e.g., managing Kubernetes deployments, grading assignments on Canvas, syncing product inventories across databases) are substantially different from Corecraft’s customer support domain, suggesting that the trained model acquired broadly applicable agentic skills rather than domain-specific shortcuts.
@@ 7 Discussion: Environment Design and Its Implications for Agent Training
The production agent deployment landscape shows that reliability remains the primary barrier to real-world adoption (Pan et al., 2025). Our results suggest that the design properties of training environments play an important role in addressing this challenge.
Many existing agent benchmarks rely on synthetic data generation to achieve scale (Qin et al., 2024). While synthetic approaches enable large task counts, they may introduce distributional artifacts that limit transfer. The Corecraft environment takes a different approach, grounding tasks in realistic professional workflows through three properties that we believe contribute to the observed generalization. First, Corecraft features diverse task types spanning information retrieval, communication, reasoning, and multi-step workflows, requiring agents to develop a broad repertoire of skills rather than narrow heuristics. This diversity of difficulty levels within the training distribution relates to insights from curriculum learning research (Narvekar et al., 2020), which has shown that exposure to varied task difficulties can improve learning outcomes in RL domains. Second, expert-authored rubrics decompose success into verifiable sub-criteria, providing a richer reward signal than binary pass/fail. This decomposition can be viewed as a form of reward shaping (Ng et al., 1999), where intermediate criteria guide the policy toward desired behaviors. The rubric-based approach shares a common insight with recent work on rubric-based training for instruction following (He et al., 2025): decomposing evaluation into fine-grained criteria enables more effective RL training. While that work demonstrated the approach for instruction following, our results extend the paradigm to multi-step agentic tasks with tool use. Third, tasks are designed to mirror authentic workplace challenges, encouraging models to learn professional patterns (e.g., structured communication, constraint-aware data retrieval) that apply across domains.
The transfer to $\tau^{2}$-Bench Retail is particularly encouraging given that customer service represents one of the most common deployment domains for production agents. Pan et al. (2025) report that finance, technology, and corporate services represent the highest-concentration deployment domains, with customer support among the most frequent application areas. Improvements in this domain could have immediate practical impact.
Our results suggest that environment quality, diversity, and realism play a meaningful role in producing transferable capabilities. A promising future direction would be to combine expert-designed core scenarios with synthetic variations to optimize for both quality and coverage.
@@ 7.1 Future Work
We evaluated only GLM 4.6. Testing whether similar gains emerge across different model families and scales would further validate that environment quality, rather than model-specific interactions, drives improvement.
@@ 8 Conclusion
We presented evidence that high-fidelity RL environments enable generalizable agent training. Using Corecraft, a realistic customer support environment, we showed that single-epoch GRPO training produces 11.39 percentage point improvement on held-out tasks and transfers to out-of-distribution benchmarks (+4.5% BFCL Parallel, +7.4% $\tau^{2}$-Bench Retail, +6.2% Toolathlon).
@@
9 Broader Impact
This work contributes to the development of more capable AI agents for enterprise settings. We discuss some of these implications below.
Positive Implications.
Improved agent reliability could reduce errors in customer service and other enterprise workflows, benefiting both businesses and customers. The transfer learning results suggest that high-quality training environments may be more efficient than scaling data quantity alone, potentially reducing computational resources required for agent development. Additionally, our emphasis on expert-designed tasks and rubrics provides a methodology for incorporating domain expertise into AI training, which could help ensure agents learn appropriate professional practices.
Mitigations.
The rubric-based evaluation methodology we employ provides transparency into agent failures, enabling targeted improvements and informed deployment decisions. We encourage practitioners to maintain human oversight for high-stakes decisions and to use evaluation frameworks that surface important limitations.
@@ 10 Acknowledgements
We thank the Surge AI workforce for their essential contributions to this research. Domain experts designed workplace tasks with complexity and realistic edge cases that reflect actual operational challenges. Annotators populated the Corecraft environment with coherent entities, relationships, and data that approximate real-world e-commerce systems. Expert evaluators assessed model trajectories and identified failure patterns, enabling the systematic analysis presented in this work. We thank Ian Niebres for setting up the CoreCraft environment and John Stacy for conducting thorough evaluations and analyses on external benchmarks.

### v2

# EnterpriseBench Corecraft: Training Generalizable Agents on High-Fidelity RL Environments

@@ Abstract
We show that training AI agents on high-fidelity reinforcement learning environments produces capabilities that generalize beyond the training distribution. We introduce Corecraft, the first environment in EnterpriseBench, Surge AI’s suite of agentic RL environments. Corecraft is a fully operational enterprise simulation of a customer support organization, comprising over 2,500 entities across 14 entity types with 23 unique tools, designed to measure whether AI agents can perform the multi-step, domain-specific work that real jobs demand. Frontier models such as GPT-5.2 and Claude Opus 4.6 solve fewer than 30% of tasks when all expert-authored rubric criteria must be satisfied. Using this environment, we train GLM 4.6 with Group Relative Policy Optimization (GRPO) and adaptive clipping. After a single epoch of training, the model improves from 25.37% to 36.76% task pass rate on held-out evaluation tasks. More importantly, these gains transfer to out-of-distribution benchmarks: +4.5% on BFCL Parallel, +7.4% on $\tau^{2}$-Bench Retail, and +6.8% on Tool Decathlon (Pass@1). We believe three environment properties are consistent with the observed transfer: task-centric world building that optimizes for diverse, challenging tasks; expert-authored rubrics enabling reliable reward computation; and enterprise workflows that reflect realistic professional patterns. Our results suggest that environment quality, diversity, and realism are key factors enabling generalizable agent capabilities.
@@ 1 Introduction
  • We show transfer to out-of-distribution benchmarks: +4.5% on BFCL Parallel function calling, +7.4% on $\tau^{2}$-Bench Retail customer service, and +6.8% on Tool Decathlon long-horizon tool use.
@@ 2.1 Agent Benchmarks and Evaluation
Customer service evaluation has been addressed by $\tau$-bench (Yao et al., 2024) and its successor $\tau^{2}$-bench (Barres et al., 2025), which introduced the passk metric measuring reliability across multiple attempts. Tool Decathlon or Toolathlon (Li et al., 2025) benchmarks agents on 108 diverse, long-horizon tasks spanning 32 software applications and 604 tools, with execution-based evaluation. AgentBoard (Ma et al., 2024) provides multi-turn evaluation with subgoal progress tracking across diverse agentic tasks.
@@ 4 Frontier Model Performance on Corecraft
To characterize the difficulty of Corecraft and establish the headroom available for training, we evaluated frontier models on the current Corecraft benchmark task set. Table 1 reports task pass rates, where a task is counted as passed only when all rubric criteria are satisfied across trajectories.
@@
Table 1: Frontier model performance on the current Corecraft benchmark task set. Pass rate requires all rubric criteria to be satisfied. Even the best model (GPT-5.2 High) solves fewer than 30% of tasks.
These results reveal substantial headroom. Even GPT-5.2, the strongest model on this set, achieves only 29.20% when all rubric criteria must be satisfied. This shows that Corecraft presents challenging tasks that expose capability limitations even in the most advanced models, providing a good training signal for RL.
@@ 5.1 Base Model and Data
We use GLM 4.6 as our base model, a 357B parameter mixture-of-experts architecture with 32B active parameters. We used a random train-test split from the Corecraft environment with 1000 tasks used for training and 150 tasks in the held-out evaluation test.
@@ 5.4 Reward Computation
where $C$ is the set of rubric criteria. This provides a dense reward signal while maintaining interpretable task-level metrics.
@@ 6.1 In-Distribution Performance
Table 3 presents task pass rates on the held-out Corecraft evaluation set used during training. [^1] Since training was conducted, new tasks have been added to Corecraft benchmark, resulting in the Corecraft benchmark set reported in Section 4. The results in Table 3 reflect the task split available at training time. A task is counted as passed only when all rubric criteria are satisfied, making this a strict measure of end-to-end task completion. After one epoch of training, GLM 4.6 improves from 25.37% to 36.76%, a gain of 11.39 percentage points.
@@ 6.3 Out-of-Distribution Generalization
Table 4 presents results on benchmarks outside the Corecraft training distribution. The trained model shows consistent improvements across long-horizon tool-use benchmarks.
@@
Toolathlon (Li et al., 2025) evaluates language agents on 108 diverse, long-horizon tasks spanning 32 software applications and 604 tools, requiring an average of approximately 20 interaction turns per task. The trained model improves from 18.8% to 25.6% (+6.8%), placing it between Claude Haiku 4.5 and Grok-4 in our evaluations. This is a particularly notable transfer result because Toolathlon tasks (e.g., managing Kubernetes deployments, grading assignments on Canvas, syncing product inventories across databases) are substantially different from Corecraft’s customer support domain, suggesting that the trained model acquired broadly applicable agentic skills rather than domain-specific shortcuts.
@@ 7 Discussion: Environment Design and Its Implications for Agent Training
The current production agent deployment landscape shows that reliability remains the main barrier to real-world adoption (Pan et al., 2025). Our results suggest that the design properties of training environments play an important role in addressing this challenge.
Many existing agent benchmarks rely on synthetic data generation to achieve scale (Qin et al., 2024). While synthetic approaches enable large task counts, they may introduce distributional artifacts that limit transfer. The Corecraft environment takes a different approach, grounding tasks in realistic professional workflows through three properties that we believe contribute to the observed generalization. First, Corecraft features diverse task types spanning information retrieval, communication, reasoning, and multi-step workflows, requiring agents to develop a broad repertoire of skills rather than narrow heuristics. This diversity of difficulty levels within the training distribution relates to insights from curriculum learning research (Narvekar et al., 2020), which has shown that exposure to varied task difficulties can improve learning outcomes in RL domains. Second, expert-authored rubrics decompose success into verifiable sub-criteria, providing a richer reward signal. Third, tasks are designed to mirror authentic workplace challenges, encouraging models to learn professional patterns (e.g., structured communication, constraint-aware data retrieval) that apply across domains.
Our results suggest that environment quality, diversity, and realism play a meaningful role in producing transferable capabilities.
@@ 7.1 Future Work
We trained and evaluated GLM 4.6. Testing whether similar gains emerge across different model families and scales would further validate that environment quality, rather than model-specific interactions, drives improvement.
@@ 8 Conclusion
We presented evidence that high-fidelity RL environments enable generalizable agent training. Using Corecraft, a realistic customer support environment, we showed that single-epoch GRPO training produces 11.39 percentage point improvement on held-out tasks and transfers to out-of-distribution benchmarks (+4.5% BFCL Parallel, +7.4% $\tau^{2}$-Bench Retail, +6.8% Toolathlon).
@@ 9 Acknowledgements
We thank the Surge AI workforce for their essential contributions to this research. Domain experts designed workplace tasks with complexity and realistic edge cases that reflect actual operational challenges. Annotators populated the Corecraft environment with coherent entities, relationships, and data that approximate real-world e-commerce systems. Expert evaluators assessed model trajectories and identified failure patterns, enabling the systematic analysis presented in this work. We thank Ian Niebres for setting up the Corecraft environment, Tim Bauman for leading Corecraft engineering and John Stacy for conducting evaluations and analyses on external benchmarks.